OpenStudy (anonymous):

Solve the following system of equations. x + y – 4z = –7 2x + 3y – 3z = –6 2x + 2y – z = 0

OpenStudy (watchmath):

-2*first + second:\(y+5z=8\) -2*first+third: \(7z=14\) Hence \(z=2\) \(y+10=8\) \(y=-2\) \(x-2-8=-7\) \(x=3\)